Show that the function defined as 𝑓(π‘₯) = { sin 1 π‘₯ , π‘₯ β‰  0 0, π‘₯ = 0 obeys the intermediate value theorem


1
Expert's answer
2021-12-16T12:15:21-0500

The function "f(x)" is continuous on "(-\infin, 0)\cup(0, \infin)."

So ifΒ "a" andΒ "b" have the same sign then such "c" Β exists by the Intermediate Value Theorem (IVT).

If "a\leq0\leq b." Β Assume "b>0."

Let "n" be a natural number such that "2/b <(2n+1)\pi" so that the interval "J=[\dfrac{2}{(2n+3)\pi},\dfrac{2}{(2n+1)\pi}]" is contained in "[0,b]."

The function "f(x) =\sin (1/x)" is continuous on "J" and takes on the values "1" and "\u22121" at the endpoints of "J."

Since "\u22121 \u2264 f(x) \u2264 1," the Intermediate Value Theorem applied to "f" on "J" implies that given any "y" such that "\u22121 < y < 1," there is "c" in "J" such that "f(c) = y."

The case "a<0" is similar.